OpenStudy (anonymous):

is The following is an arithmetic sequence. 2, 5, 8, 10,…

OpenStudy (anonymous):

looks like \(5-2=3\) and \(8-5=3\) but \(10-8=2\) so not, the differences are not the same

OpenStudy (anonymous):

@satellite73 so no? thanks!